Use the given probability distribution table to answer each item.


X 0 1 2 3 4 5


P(X) 0.10 0.24 0.08 0.20 0.21 0.17



1. P(2)=


2.P(greater than 0)=


3.P(less than 5)=


4.P(even)=


5.P(odd)=


6.P(between 0 and 5)=


7.P(greater than 2)=


8.P(less than 3)=


9.P(3 and 4)=


10.<P(X)=


1
Expert's answer
2022-02-22T00:51:00-0500

Solution:

X 0 1 2 3 4 5

P(X) 0.10 0.24 0.08 0.20 0.21 0.17

1. P(2)=0.08

2.P(greater than 0)="P(X>0)=1-P(X\le0)=1-P(X=0)=1-0.10=0.9"

3.P(less than 5)="P(X<5)=1-P(X\ge5)=1-P(X=5)=1-0.17=0.83"

4.P(even)=P(0)+P(2)+P(4)=0.10+0.08+0.21=0.39

5.P(odd)=1-P(even)=1-0.39=0.61

6.P(between 0 and 5)=1-P(0)-P(5)=1-0.10-0.17=1-0.27=0.73

7.P(greater than 2)=P(3)+P(4)+P(5)=0.20+0.21+0.17=0.58

8.P(less than 3)=P(0)+P(1)+P(2)=0.10+0.24+0.08=0.42

9.P(3 and 4)=P(3)+P(4)=0.20+0.21=0.41

10.<P(X)= Incorrectly written